# Generative Adversarial Gumbel MCTS for Abstract Visual Composition Generation

### v1

@@ 1 Introduction
Abstract visual concepts, those defined by structural relations among primitives rather than by rich texture or appearance, underlie many forms of human reasoning, communication and design. For example, architectural blueprints, engineering schematics, iconography and signage compose meaning from simplified geometric forms across contexts. We therefore foreground visual composition: meaning emerges from how simple parts are arranged into structure. We call a visual concept abstract when its identity is specified primarily by structural regularities over simple primitives (shape, part relations, closure, symmetry), consistent with structural accounts and Gestalt principles of perceptual organisation [biederman1987rbc, wertheimer1923gestalt]. In this paper, we study composition explicitly by assembling fixed geometric pieces into target silhouettes or layouts from weak language cues (e.g., “bird,” “boat”), e.g., Tangram Assembly in Fig 1.
@@
Composing with fixed pieces under geometric constraints differs fundamentally from naturalistic synthesis. First, feasibility is discrete and brittle: a single overlap or misalignment invalidates a composition; the feasible set is sparse and highly non-convex. Second, search is combinatorial: at each step, the system must choose which piece, where, and how to orient it while preserving future flexibility. Related packing/cutting problems are NP-hard even in simplified rectangular settings, underscoring intrinsic difficulty [garey1979np, iori2021twodpack]. Third, semantics are underspecified: textual prompts imply global shape regularities without prescribing unique layouts. Data-driven generators trained to match pixel statistics (e.g., diffusion, latent diffusion) excel at appearance but struggle to guarantee hard constraints or discover valid layouts from limited supervision [ho2020ddpm, rombach2022ldm]. Object-centric auto-regressive transformer approaches argue for compositional structure, but still lack exact constraint enforcement during synthesis [locatello2020slot].
@@
To address these challenges, we propose an AlphaGo-like framework that integrates hard constraints into Monte Carlo Tree Search (MCTS) and employs a vision-language model to assess whether a generated configuration matches the textual goal. The key assumption is that validating an abstract visual composition is easier than learning to compose it directly. Prior work shows that fine-tuned vision-language models can reach human-level accuracy in recognising abstract tangram shapes [ji2022abstract], suggesting their suitability as reward models for evaluating candidate solutions.
@@
Our method adopts an AlphaGo-style algorithm [silver2018alphazero], where hard constraints are embedded in the transition model to prune infeasible actions. A fine-tuned CLIP model [radford2021clip] serves as the reward function, measuring the semantic alignment between the final assembly and the goal description. Because such scorers may be permissive or inconsistent on out-of-distribution samples, we introduce adversarial reward refinement to better distinguish valid from near-miss compositions. The policy guiding MCTS is trained using Proximal Policy Optimisation [schulman2017proximal] (PPO) with the refined reward, and the resulting policy is further used to generate hard negative examples for continued adversarial reward refinement. This framework enhances sample efficiency through constraint-based pruning and improves semantic faithfulness by leveraging language-aligned visual features.
@@
We evaluate on Tangram assembly: seven fixed pieces must depict a textual concept; this probes joint semantic and geometric competence. We found that methods without search are limited in generating semantically aligned abstract visual concepts. To better understand the issues, we relax the problem into a rectangle composition task, where a composition is semantically valid if all pieces lie within a bounded area of width and height specified by the text. We found that the diffusion model and PPO achieved great performance in generation with weak geometrical constraints, but performed poorly with tight geometrical constraints. Across both, the proposed search-based generator achieves higher geometric validity and semantic agreement than diffusion and autoregressive baselines, with gains increasing under tighter constraints [iori2021twodpack].
@@ 2 Related work
Abstract visual reasoning. Abstract visual concepts are intangible ideas or notions that lack a physical form or concrete attributes. They cannot be directly perceived and typically denote qualities, relationships, or processes [borghi2022abstract]. Those concepts have long been used to explore how people collaborate in communication when there are no established naming conventions, dating back to [krauss1964changes]. Tangrams were introduced as a research tool by [clark1986referring]. Although these figures are composed of the same seven basic shapes, they inspire a diverse range of figurative descriptions, reflecting different ways of conceptualising them [horton2002speakers, duff2006development, holler2011co, castillo2019interaction, bangerter2020lexical]. A recent study [ji2022abstract] focuses on the tangram reasoning for classification purposes using modern vision-language models. Those models process multimodal data through joint encoding, which captures richer interactions [lu2019vilbert, chen2020cops, kim2021vilt], or separate encoding, which is more flexible [radford2021learning, jia2021scaling].
@@
For generating/visualising abstract visual concepts, there are a much smaller number of studies. Some research works use a diffusion model to generate images with abstract style or vector format (e.g., SVG) given language descriptions [frans2022clipdraw, schaldenbrand2022styleclipdraw, xing2023diffsketcher, wu2023iconshop, jain2023vectorfusion] or photo-realistic images [ma2022towards, li2020differentiable]. Some studies [liao2024text, liu2022opal, xiao2015vismantic, cunha2020visual, ge2021visual, sun2025creative] also focus on visualising highly abstract concepts with metaphor, which is highly subjective and does not focus on generating abstract representations of concrete objects. These works treat the language or image input as soft constraints, and many are very successful. However, neither of them had the hard geometrical constraints for their generation tasks.
@@
Constrained Generation. Constrained generation is increasingly important for outputs that must satisfy geometric or physical rules. Methods incorporate constraints during training or inference, including geometric-loss GANs [zeng2021enforcing] and factor-graph layouts [dupty2024constrained]. In CAD and mechanical design, models such as SketchGen [para2021sketchgen] emit primitives plus constraints for downstream solvers; dual-discriminator GANs [almasri2020shape] and parameter-efficient models [padula2024generative] target mechanical and geometric validity. Together, these works integrate domain constraints into generative pipelines, improving reliability.
@@
Diffusion models show promise but struggle to enforce hard geometric rules: denoising approximates constraints without guarantees. Post-hoc projections (mirror mapping [liu2023mirror], optimisation-based sampling [christopher2024constrained]) remain approximate and can fail under novel configurations; training-free fixes [zampini2025training] add heuristics without verification. Without explicit geometric representations, constraints become negotiable preferences—unacceptable in precision domains (e.g., mechanical design). Recent MCTS-guided diffusion planning [yoon2025monte] uses rewards, but its action space still depends on diffusion via two meta actions (guide/no_guide), keeping hard constraints implicit and unenforced.
@@ 5.1 Experiment setup
  • Classifier-Guided Diffusion model. This baseline is to use the classifier-guided diffusion model [dhariwal2021diffusion] to generate the configurations of the tangram pieces. We use the same neural network architecture as our method, but we replace the MCTS with the diffusion model. To make the comparison fair, we also design a differentiable classifier that encodes the geometrical constraints to guide the diffusion model to generate the configurations of tangram pieces. For more details, please refer to the Appendix D. We provide the results with different classifier guidance scales $\beta$: stronger scales $\beta$ mean greater classifier guidance.
@@ Appendix A Preliminary: Gumbel Muzero
Gumbel Muzero [danihelka2022policy] is a variant of the AlphaZero [silver2017mastering] algorithm with a major modification in the MCTS algorithm. Instead of using the upper confidence bound, Gumbel Muzero uses sequential halving, which minimises simple regrets, for the root action selection. As such, it is guaranteed to have policy improvement even with the minimal number of simulations in MCTS during the training.
@@
Gumbel Muzero uses the Gumbel-Max trick [gumbel1954statistical, yellott1977relationship, jang2016categorical, kool2019stochastic] to sample $n$ actions without replacement:
@@
Sequential Halving [karnin2013almost] is a bandit algorithm for simple regret minimisation. It aims at maximising the $Q$-value of the last or the finally selected action in simulations, which differs from the cumulative regrets that maximise the $Q$ from all $n$ simulations [tolpin2012mcts].
@@ D.1 GAG Muzero
  • Policy/Value Network: A Decision Transformer [chen2021decision] backbone is used, with BERT-based text encoding and a vision transformer (ViT) for image processing. The network outputs action logits and value estimates.
  • Reward Model: A CLIP-based [radford2021learning] reward model computes the alignment between the assembled image and the instruction, using contrastive and pairwise preference losses.
@@
  • Self-Play: The agent generates training data via self-play, using a Gumbel MuZero Monte Carlo Tree Search (MCTS) [schrittwieser2020mastering] for action selection. The search incorporates a recurrent function that simulates environment transitions and model predictions.

### v2

@@ 1 Introduction
Abstract visual concepts, those defined by structural relations among primitives rather than by rich texture or appearance, underlie many forms of human reasoning, communication and design. For example, architectural blueprints, engineering schematics, iconography and signage compose meaning from simplified geometric forms across contexts. We therefore foreground visual composition: meaning emerges from how simple parts are arranged into structure. We call a visual concept abstract when its identity is specified primarily by structural regularities over simple primitives (shape, part relations, closure, symmetry), consistent with structural accounts and Gestalt principles of perceptual organisation [3, 50]. In this paper, we study composition explicitly by assembling fixed geometric pieces into target silhouettes or layouts from weak language cues (e.g., “bird,” “boat”), e.g., Tangram Assembly in Fig 1.
@@
Composing with fixed pieces under geometric constraints differs fundamentally from naturalistic synthesis. First, feasibility is discrete and brittle: a single overlap or misalignment invalidates a composition; the feasible set is sparse and highly non-convex. Second, search is combinatorial: at each step, the system must choose which piece, where, and how to orient it while preserving future flexibility. Related packing/cutting problems are NP-hard even in simplified rectangular settings, underscoring intrinsic difficulty [16, 22]. Third, semantics are underspecified: textual prompts imply global shape regularities without prescribing unique layouts. Data-driven generators trained to match pixel statistics (e.g., diffusion, latent diffusion) excel at appearance but struggle to guarantee hard constraints or discover valid layouts from limited supervision [19, 42]. Object-centric auto-regressive transformer approaches argue for compositional structure, but still lack exact constraint enforcement during synthesis [35].
@@
To address these challenges, we propose an AlphaGo-like framework that integrates hard constraints into Monte Carlo Tree Search (MCTS) and employs a vision-language model to assess whether a generated configuration matches the textual goal. The key assumption is that validating an abstract visual composition is easier than learning to compose it directly. Prior work shows that fine-tuned vision-language models can reach human-level accuracy in recognising abstract tangram shapes [25], suggesting their suitability as reward models for evaluating candidate solutions.
@@
Our method adopts an AlphaGo-style algorithm [46], where hard constraints are embedded in the transition model to prune infeasible actions. A fine-tuned CLIP model [40] serves as the reward function, measuring the semantic alignment between the final assembly and the goal description. Because such scorers may be permissive or inconsistent on out-of-distribution samples, we introduce adversarial reward refinement to better distinguish valid from near-miss compositions. The policy guiding MCTS is trained using Proximal Policy Optimisation [45] (PPO) with the refined reward, and the resulting policy is further used to generate hard negative examples for continued adversarial reward refinement. This framework enhances sample efficiency through constraint-based pruning and improves semantic faithfulness by leveraging language-aligned visual features.
@@
We evaluate on Tangram assembly: seven fixed pieces must depict a textual concept; this probes joint semantic and geometric competence. We found that methods without search are limited in generating semantically aligned abstract visual concepts. To better understand the issues, we relax the problem into a rectangle composition task, where a composition is semantically valid if all pieces lie within a bounded area of width and height specified by the text. We found that the diffusion model and PPO achieved great performance in generation with weak geometrical constraints, but performed poorly with tight geometrical constraints. Across both, the proposed search-based generator achieves higher geometric validity and semantic agreement than diffusion and autoregressive baselines, with gains increasing under tighter constraints [22].
@@ 2 Related work
Abstract visual reasoning. Abstract visual concepts are intangible ideas or notions that lack a physical form or concrete attributes. They cannot be directly perceived and typically denote qualities, relationships, or processes [4]. Those concepts have long been used to explore how people collaborate in communication when there are no established naming conventions, dating back to [30]. Tangrams were introduced as a research tool by [9]. Although these figures are composed of the same seven basic shapes, they inspire a diverse range of figurative descriptions, reflecting different ways of conceptualising them [21, 13, 20, 5, 2]. A recent study [25] focuses on the tangram reasoning for classification purposes using modern vision-language models. Those models process multimodal data through joint encoding, which captures richer interactions [36, 7, 28], or separate encoding, which is more flexible [41, 26].
@@
For generating/visualising abstract visual concepts, there are a much smaller number of studies. Some research works use a diffusion model to generate images with abstract style or vector format (e.g., SVG) given language descriptions [15, 43, 53, 51, 23] or photo-realistic images [37, 31]. Some studies [32, 34, 52, 10, 17, 48] also focus on visualising highly abstract concepts with metaphor, which is highly subjective and does not focus on generating abstract representations of concrete objects. These works treat the language or image input as soft constraints, and many are very successful. However, neither of them had the hard geometrical constraints for their generation tasks.
@@
Constrained Generation. Constrained generation is increasingly important for outputs that must satisfy geometric or physical rules. Methods incorporate constraints during training or inference, including geometric-loss GANs [57] and factor-graph layouts [14]. In CAD and mechanical design, models such as SketchGen [39] emit primitives plus constraints for downstream solvers; dual-discriminator GANs [1] and parameter-efficient models [38] target mechanical and geometric validity. Together, these works integrate domain constraints into generative pipelines, improving reliability.
@@
Diffusion models show promise but struggle to enforce hard geometric rules: denoising approximates constraints without guarantees. Post-hoc projections (mirror mapping [33], optimisation-based sampling [8]) remain approximate and can fail under novel configurations; training-free fixes [56] add heuristics without verification. Without explicit geometric representations, constraints become negotiable preferences—unacceptable in precision domains (e.g., mechanical design). Recent MCTS-guided diffusion planning [55] uses rewards, but its action space still depends on diffusion via two meta actions (guide/no_guide), keeping hard constraints implicit and unenforced.
@@ 5.1 Experiment setup
  • Classifier-Guided Diffusion model. This baseline is to use the classifier-guided diffusion model [12] to generate the configurations of the tangram pieces. We use the same neural network architecture as our method, but we replace the MCTS with the diffusion model. To make the comparison fair, we also design a differentiable classifier that encodes the geometrical constraints to guide the diffusion model to generate the configurations of tangram pieces. For more details, please refer to the Appendix D. We provide the results with different classifier guidance scales $\beta$: stronger scales $\beta$ mean greater classifier guidance.
@@ Appendix A Preliminary: Gumbel Muzero
Gumbel Muzero [11] is a variant of the AlphaZero [47] algorithm with a major modification in the MCTS algorithm. Instead of using the upper confidence bound, Gumbel Muzero uses sequential halving, which minimises simple regrets, for the root action selection. As such, it is guaranteed to have policy improvement even with the minimal number of simulations in MCTS during the training.
@@
Gumbel Muzero uses the Gumbel-Max trick [18, 54, 24, 29] to sample $n$ actions without replacement:
@@
Sequential Halving [27] is a bandit algorithm for simple regret minimisation. It aims at maximising the $Q$-value of the last or the finally selected action in simulations, which differs from the cumulative regrets that maximise the $Q$ from all $n$ simulations [49].
@@ D.1 GAG Muzero
  • Policy/Value Network: A Decision Transformer [6] backbone is used, with BERT-based text encoding and a vision transformer (ViT) for image processing. The network outputs action logits and value estimates.
  • Reward Model: A CLIP-based [41] reward model computes the alignment between the assembled image and the instruction, using contrastive and pairwise preference losses.
@@
  • Self-Play: The agent generates training data via self-play, using a Gumbel MuZero Monte Carlo Tree Search (MCTS) [44] for action selection. The search incorporates a recurrent function that simulates environment transitions and model predictions.